If my length is 1 and 6 over 7 and my width is 7/8 what is my area and perimeter

Answer by checkley79(3341)

AREA=LW
(1 6/7)(7/8)
[(7+6)/7](7/8)
(13/7)(7/8)
13/8=1.625 ANS. FOR THE AREA.
PERIMETER=2L+2W
P=2(13/7)+2(7/8)
P=26/7+14/8
P=(26*8+14*7)/7*8
P=(208+98)/56
P=306/56
P=5.46 ANS FOR THE PERIMETER.